OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

My suggest would be to try and use the first equation to solve for y, in terms of x. Then using that solution, replace y in the second equation and solve for x!

OpenStudy (shane_b):

To use substitution you just solve one equation in terms of either x or y...then plug that equation into the other original equation and solve for the missing variable. Original equations: 2x + y = 1 4x + 2y = −2 Let's go with the first equation and solve for y: 2x + y = 1 (original equation) Therefore y=1-2x Plug that equation into the the other (second original) equation: 4x + 2y = -2 (original equation) 4x + 2(1-2x)=-2 4x+2-4x=-2 Simplify to get: 2 = -2 Since that statement is wrong, there is no unique solution for the given equations.